find the slope-intercept equation for the line with the indicated slope and y-intercept slope 5/7; y-intercept (0,4)
Answer by TutorDelphia(193)

y=mx+b is slope intercept for where m=the slope and b=the intercept
In this case the slope is 5/7 so m=5/7
The intercept is 4 because when x=0 and y=4 is where the line will intercept the y axis.
y=(5/7)x+4